# RanDumb: A Simple Approach that Questions the Efficacy of Continual Representation Learning

### v1

@@ Abstract
We propose RanDumb to examine the efficacy of continual representation learning. RanDumb embeds raw pixels using a fixed random transform which approximates an RBF-Kernel, initialized before seeing any data, and learns a simple linear classifier on top. We present a surprising and consistent finding: RanDumb significantly outperforms the continually learned representations using deep networks across numerous continual learning benchmarks, demonstrating the poor performance of representation learning in these scenarios. RanDumb stores no exemplars and performs a single pass over the data, processing one sample at a time. It complements GDumb [39], operating in a low-exemplar regime where GDumb has especially poor performance. We reach the same consistent conclusions when RanDumb is extended to scenarios with pretrained models replacing the random transform with pretrained feature extractor. Our investigation is both surprising and alarming as it questions our understanding of how to effectively design and train models that require efficient continual representation learning, and necessitates a principled reinvestigation of the widely explored problem formulation itself. Our code is available here.
@@ 1 Introduction
Continual learning is a specialized form of supervised learning, characterized by sequentially arriving tasks, coupled with additional computational and memory constraints [24, 27, 44, 49, 39] (see Verwimp et al. [51] for a survey of applications). A notable implication of this setup is the inherent performance gap between continual learning and traditional supervised learning. To elucidate this further, let us consider the achievable risk under continual learning given a dataset $\mathcal{D}$:
$\mathcal{R}_{\texttt{CL}}(\mathcal{D})\geq R_{\texttt{Joint}}(\mathcal{D})+E_{\texttt{CL}}(\mathcal{D})$ (1)
Here, $R_{\texttt{Joint}}$ denotes the risk achievable with supervised learning without any constraints, while $E_{\texttt{CL}}$ represents the gap arising from transitioning from fully supervised learning to continual learning which imposes additional constraints such as: resource-constraint (e.g., single-epoch), distribution shift, and predefined task ordering with limited access to data. Because of these additional constraints imposed by continual learning, the risk is non-decreasing and $E_{\texttt{CL}}$ is non-negative. Minimizing $E_{\texttt{CL}}$ has been the primary focus of the continual learning literature [16, 29, 37, 41].
Figure 1: Mechanism of RanDumb. Project raw pixels to a high dimensional space using random fourier projections ($\varphi$), then decorrelate the features using Mahalanobis distance [31] and classify with the nearest class mean. The online update only involves updating a single covariance matrix and class-means.
In this work, we discover that the fundamental aspect of continual representation learning (CRL) is often overlooked in the continual learning literature, as most methods merge continual classifier and representation learning. We dissect these two aspects and study the effect of representation learning on $E_{\texttt{CL}}$. The central question of our work is trying to answer:
Are current continual learning setups overly constrained for effective continual representation learning?
Specifically, consider a continual representation learning algorithm $\mathtt{CRL}$ which learns a representation from the continual dataset $\mathcal{D}$. Is this representation better than a random (ie. non-learned) representation RandRep? More formally, is the following quantity positive or negative?
$RepGain(\mathcal{D})=E_{\texttt{CL}}(\mathcal{D},\mathtt{RandRep})-E_{\texttt{CL}}(\mathcal{D},\mathtt{CRL}(\mathcal{D}))$
Here, the random representation functions are defined at initialization without seeing any data or using domain-specific priors. Note that one expects the impact of continual representation learning ($RepGain(\mathcal{D})$) to be positive, as continual representation learning approaches should improve performance over even the best random representation.
Our primary contribution is empirically demonstrating that currently $RepGain(\mathcal{D})<0$, i.e. there exists random representations which outperform state-of-the-art representations learned by continual learning algorithms using deep networks. We show this by introducing a straightforward baseline, which we name RanDumb, standing for Random representation function and a Dumb linear classifier. Despite replacing the deep embedder with a random function, we see from Table 1 (top) that RanDumb outperforms the current state-of-the-art methods being exemplar-free over challenging large task scenarios given in [59]. We extend the large task scenario to methods with pretrained feature extractors [64]. In RanDumb, we replace our random projection with the pretrained feature extractor here, simply learning a linear classifier. We show results in Table 2 (top), showing that RanDumb outperforms the best methods in this scenario as well.
Table 1: Online Continual Learning. Performance comparison of RanDumb on the PEC setup [59] and VAE-GC [50]. Setup and numbers borrowed from PEC [59]. RanDumb outperforms the best exemplar-free OCL method. Comparing representation learning with no representation learning, a fixed, random function outperforms training a deep network continually, i.e. $RepGain(\mathcal{D})<0$. Furthermore, RanDumb nearly matches performance of one-pass joint, demonstrating the inefficacy of current benchmarks.
Method MNIST CIFAR10 CIFAR100 m-ImNet
@@
Only Ablating Representation Learning: Random v/s Deep Embedding
@@
Table 2: Offline Continual Learning. Performance comparison with ImageNet21K ViT-B16 model using 2 init classes and 1 new class per task. RanDumb here assumes the features as input pixels. SLCA [62] and RanPAC [30] results reproduced from original codebase. Comparing representation learning with no representation learning, a fixed, pretrained feature extractor outperforms further finetuning a deep network continually, i.e. $RepGain(\mathcal{D})<0$. RanDumb nearly matches performance of joint, demonstrating the inefficacy of current benchmarks. *here representation learning collapses, accuracy reported without PETL. ${}^{+}$here representation learning and random projection both collapse, reporting last stable accuracy.
@@
Best (RanPAC) 89.6 ${}^{+}$ 67.3 ${}^{+}$ 77.9* ${}^{+}$ $*{}^{*}$
RanDumb (Ours) 86.8 42.2 64.9 88.5 75.3 92.3 69.1
Improvement -2.9 +15.4 -2.4 +1.3 -2.6 +3.7 +15.4
Only Ablating Representation Learning: No Finetuning v/s Full Finetuning
@@
RanDumb (Ours) 86.8 42.2 64.9 88.5 75.3 92.3 69.1
Improvement +0.0 - +10.7 +6.4 - - +50.9
@@
RanDumb (Ours) 86.8 42.2 64.9 88.5 75.3 92.3 69.1
Gap Covered. (%) 93% 60% 75% 97% 92% 97% 80%
Figure 2: Mechanism of RanDumb. RanDumb projects the input datapoints to a high-dimensional space to create a clearer separation between classes. Subsequently, it corrects the anisotropy across different feature dimensions, scaling them to be unit variance each. This allows a nearest class mean classifier to accurately rely on cosine similarity. The figure is adapted from Pilario et al. [38].
1.1 RanDumb: Construction and Performance
@@
Mechanism. Our mechanism is illustrated in Figure 1. RanDumb first projects input pixels into a high-dimensional space using a fixed kernel based on random Fourier basis, which is a low-rank data-independent approximation of the RBF Kernel [42]. Then, we use a simple linear classifier which first normalizes distances across different feature dimensions (anisotropy) with Mahalanobis distance [31] and then uses nearest class means for classification [33]. In scenarios with pretrained feature extractors, we use the fixed pretrained model as embedder and learn a linear classifier over it, similar to Hayes and Kanan [20].
@@
Properties. RanDumb needs no storage of exemplars and requires only one pass over the data in a one-sample-per-timestep fashion. Furthermore, it only requires online estimation of the covariance matrix and nearest class mean, hence is invariant to data order, representing the accuracy in the worst-case ordering.
@@
Poor Representation Learning. We first compare RanDumb with VAE-GC [50] in Table 1 (middle) and SLCA [62] in Table 2 (middle). The only distinction between them is their embedding functions: RanDumb uses a fixed function (random/pretrained network) for the embedder, whereas VAE-GC and SLCA further continually trained deep networks, using them as the embedder. Astonishingly, RanDumb consistently surpasses VAE-GC and SLCA baselines by wide margins of 5-15%. This shows that state-of-the-art online continual learning algorithms fail to learn effective representations in a single pass across standard online continual learning datasets.
Benchmarks Over-Constrained. The reader might now ask: With better representation learning, can one outperform RanDumb by large margins in these settings? To study this, we compare the performance of RanDumb with Joint in both online and offline settings respectively in Table 1 (bottom) and Table 2 (bottom). We observe that our simple baseline RanDumb recovers 70-90% of the gap to the respective joint classifiers across both settings. Improving forgetting in continual representation learning [21, 36] offers limited scope for improvement, as current benchmarks might be too constrained to have effective representation learning at all.
In the next sections, we shall first describe our proposed method and then walk through our findings in detail.
@@ 2 RanDumb: Mechanism & Intuitions
In this section, we thoroughly examine RanDumb, focusing on its two main elements: random projection and the dumb learner. We illustrate the mechanism of RanDumb using three toy examples in Figure 2.
We first describe the dumb classifier. To classify a test sample $\bf x_{\rm test}$, we start with a simple classifier, the nearest class mean (NCM) classifier. It predicts the class among $C$ classes by highest value of the similarity function $f$ among class means $\mathbf{\mu}_{i}$:
$y_{\textrm{pred}}=\operatorname*{arg\,max}_{{i}\in\{1,\ldots,|C|\}}f({\bf x_{\textrm{test}}},\mathbf{\mu}_{i}),\text{where}$
$f({\bf x_{\textrm{test}}},\mathbf{\mu}_{i}):={\bf x_{\textrm{test}}}^{\top}{\mathbf{\mu}_{i}}$ (2)
and $\mathbf{\mu}_{i}$ are the class-means in the pixel space:
$\mathbf{\mu}_{i}=\frac{1}{|C_{i}|}\sum_{\mathbf{x}}\mathbf{x}\qquad\forall{\bf x}\in C_{i}$
RanDumb adds two additional components to this classifier: 1) Kernelization and 2) Decorrelation.
1
Kernelization: Why & How? Classes are typically not linearly separable in the pixel space, unlike in the feature space of deep models. Hence, we apply the kernel trick to embed the pixels in a better representation space, computing all distances between the data and class-means in this embedding space. This phenomena is illustrated on three toy examples to build intuitions in Figure 2 (Embed). We use an RBF-Kernel, which for two points $\mathbf{x}$ and $\mathbf{y}$ is defined as: $K_{\text{RBF}}(\mathbf{x},\mathbf{y})=\exp\left(-\gamma\|\mathbf{x}-\mathbf{y}\|^{2}\right)$ where $\gamma$ is a scaling parameter. However, calculating the RBF kernel is not possible due to the online continual learning constraints preventing computation of pairwise-distance between all points. Hence, we use a data-independent approximation, random Fourier projection $\phi(\mathbf{x})$, as given in [42]:
@@
We keep these $\omega$ bases fixed throughout online learning. Thus, we obtain our modified similarity function from Equation 2 as:
@@
$\bar{\mathbf{\mu}}_{i}=\frac{1}{|C_{i}|}\sum_{\mathbf{x}}\phi(\mathbf{x})\qquad\forall{\bf x}\in C_{i}$
2
Decorrelation: Why & How? Projected raw pixels, similar to features of deep models, have feature dimensions with different variances (anisotropic). Hence, instead of naively computing ${\bf\phi(x_{\rm test})}^{\top}{\bar{\mathbf{\mu}}_{i}}$, we further decorrelate the feature dimensions using a Mahalonobis distance with the empirical covariance matrix $\mathbf{S}$. We illustrate this phenomena as well on three toy examples in Figure 2 (Decorrelate) to build intuitions. Our similarity function finally is:
@@
Online Computation. Our random projection is computed and fixed before seeing any real data. In the continual learning process, we only update the running class mean and empirical covariance matrix, which is accurately estimable online [^1] Online prediction is possible using the Sherman–Morrison formula to update the inverse of the covariance matrix. However, we only needed to compute inverse at the last timestep for all our experiments, hence we do not experiment with this..
Note on Equivalences. For the curious reader we draw equivalences between the linear classifier used in this work and related works. These equivalences rely on the assumption that the classes are equiprobable, which is the case for most datasets here. In these cases, nearest class mean classifier with the Mahalanobis distance metric is equivalent to linear discriminant analysis classifier [32]. Hence, one could say RanDumb is equivalent to a Streaming LDA classifier with an approximate RBF Kernel. Alternatively, one could think of the decorrelation operation as explicitly decorrelating the features with ZCA whitening [7]. We hope these alternative perspectives allow for a better understanding of the RanDumb classifier.
Table 3: Benchmark (Overview). We evaluate RanDumb across a diverse set of benchmarks that progressively relax constraints compared to RanDumb. Benchmark A closely matches RanDumb with one class per timestep and no stored exemplars. Benchmark B, D, E progressively relax the constraints on exemplars and classes per timestep. Benchmark C and E remove the online constraint by allowing unrestricted training and sample access within a task without exemplar-storage of past tasks. Benchmark F allows using large pretrained models, modified by us with one class per task, inspired by challenges detailed in Benchmark A.
@@
Table 4: Benchmark A (Ref: Table 1 from PEC [59]). We perform comprehensive comparisons of RanDumb with popular online continual learning approaches in a 1-class per task setting referred as ‘Dataset (num_tasks/1)’. This setting is a challenging setting for OCL methods except GDumb. We observe that RanDumb outperforms all approaches across all datasets by 2-6% margins, with an exception of latest work PEC [59] on CIFAR10.
Method Memory MNIST (10/1) CIFAR-10 (10/1) CIFAR-100 (100/1) miniImageNet (100/1)
@@
Rehearsal-Based
ER [13] (ICML-W ’19) 500 84.4$\pm$ 0.3 40.6$\pm$ 1.1 12.5$\pm$ 0.3 5.7$\pm$ 0.2
A-GEM [12] (ICLR ’19) 500 59.8$\pm$ 0.8 10.2$\pm$ 0.1 1.0$\pm$ 0.0 1.1$\pm$ 0.1
iCaRL [44] (CVPR ’17) 500 83.1$\pm$ 0.3 37.8$\pm$ 0.4 5.7$\pm$ 0.1 7.5$\pm$ 0.1
BiC [56] (CVPR ’19) 500 86.0$\pm$ 0.4 35.9$\pm$ 0.4 6.4$\pm$ 0.3 1.5$\pm$ 0.1
ER-ACE [10] (ICLR ’22) 500 87.8$\pm$0.2 39.9$\pm$0.5 8.2$\pm$0.2 5.7$\pm$0.2
DER [9] (NeurIPS ’20) 500 91.7$\pm$ 0.1 40.0$\pm$ 1.5 1.0$\pm$ 0.1 1.0$\pm$ 0.0
DER++ [9] (NeurIPS ’20) 500 91.9$\pm$ 0.2 35.6$\pm$ 2.4 6.2$\pm$ 0.4 1.4$\pm$ 0.1
X-DER [8] (TPAMI ’22) 500 83.0$\pm$ 0.1 43.2$\pm$ 0.5 15.6$\pm$ 0.1 8.2$\pm$ 0.4
GDumb [39] (ECCV ’20) 500 91.0$\pm$0.2 50.7$\pm$0.7 8.2$\pm$0.2 -

### v2

@@ Abstract
Continual learning has primarily focused on the issue of catastrophic forgetting and the associated stability-plasticity tradeoffs. However, little attention has been paid to the efficacy of continually learned representations, as representations are learned alongside classifiers throughout the learning process. Our primary contribution is empirically demonstrating that existing online continually trained deep networks produce inferior representations compared to a simple pre-defined random transforms. Our approach embeds raw pixels using a fixed random transform, approximating an RBF-Kernel initialized before any data is seen. We then train a simple linear classifier on top without storing any exemplars, processing one sample at a time in an online continual learning setting. This method, called RanDumb, significantly outperforms state-of-the-art continually learned representations across all standard online continual learning benchmarks. Our study reveals the significant limitations of representation learning, particularly in low-exemplar and online continual learning scenarios. Extending our investigation to popular exemplar-free scenarios with pretrained models, we find that training only a linear classifier on top of pretrained representations surpasses most continual fine-tuning and prompt-tuning strategies. Overall, our investigation challenges the prevailing assumptions about effective representation learning in online continual learning. Our code is available here.
@@ 1 Introduction
Figure 1: RanDumb projects raw pixels to a high dimensional space using random Fourier projections ($\varphi$), then decorrelate the features using Mahalanobis distance [42] and classify with the nearest class mean. The online update only involves updating a single sample covariance matrix and class-means.
Continual learning aims to develop models capable of learning from non-stationary data streams, inspired by the lifelong learning abilities exhibited by humans and the prevalence of such real-world applications (see Verwimp et al. [64] for a survey). It is characterized by sequentially arriving tasks, coupled with additional computational and memory constraints [33, 38, 54, 62, 49].
Table 1: (Left) Online Continual Learning. Performance comparison of RanDumb on the PEC setup [75] and VAE-GC [63]. Setup and numbers borrowed from PEC [75]. RanDumb outperforms the best OCL method. (Right) Offline Continual Learning. Performance comparison with ImageNet21K ViT-B16 model using 2 initial classes and 1 new class per task. RanPAC-imp is an improved version of the RanPAC code which mitigates the instability issues in RanPAC. RanDumb nearly matches performance of joint for both online and offline, demonstrating the inefficacy of current benchmarks.
Method MNIST CIFAR10 CIFAR100 m-IMN
@@
Random vs. Learned Representations
@@
Best (RanPAC-imp) 89.4 33.8 69.4 89.6 75.3 91.9 57.3
RanDumb (Ours) 86.8 42.2 64.9 88.5 75.3 92.4 67.1
Improvement -2.6 +8.4 -4.5 -1.1 +0.0 +0.5 +9.8
Random vs. Finetuned Representations
@@
RanDumb (Ours) 86.8 42.2 64.9 88.5 75.3 92.4 67.1
Improvement +0.0 - +10.7 +6.4 - - +48.9
@@
RanDumb (Ours) 86.8 42.2 64.9 88.5 75.3 92.4 67.1
Gap Covered. (%) 93% 60% 75% 97% 92% 97% 77%
Building on the foundations of supervised deep learning, the prevalent approach in continual learning has been to jointly train representations alongside classifiers. This approach follows from the assumption that learned representations are expected to outperform fixed representation functions such as kernel classifiers, as demonstrated in supervised deep learning [34, 23, 57]. However, this assumption is never validated in continual learning, with scenarios having limited updates where networks might not be trained until convergence, such as online continual learning (OCL).
In this paper, we study the efficacy of representations derived from continual learning algorithms. Surprisingly, our findings suggest that these representations might not be as beneficial as presumed. To test this, we introduce a simple baseline method named RanDumb, which combines a random representation function with a straightforward linear classifier, illustrated in Figure 1. Our empirical evaluations, summarized in Table 1 (left, top), reveal that despite replacing the representation learning with a pre-defined random representation, RanDumb surpasses current state-of-the-art across latest online continual learning benchmarks [75].
We further expand our evaluations to scenarios that use pre-trained feature extractors [67]. By substituting our random projections with these feature extractors and retaining the linear classifier, RanDumb again outperforms leading methods as shown in Table 1 (right, top).
1.1 Technical Summary: Construction of RanDumb and Empirical Findings
@@
Design. RanDumb first projects input pixels into a high-dimensional space using a fixed kernel based on random Fourier basis, which is a low-rank data-independent approximation of the RBF Kernel [52]. Then, we use a simple linear classifier which first normalizes distances across different feature dimensions (anisotropy) with Mahalanobis distance [42] and then uses nearest class means for classification [44] (mechanism illustrated in Figure 2). In scenarios with pretrained feature extractors, we use the fixed pretrained model as embedder and learn a linear classifier as described above, similar to Hayes and Kanan [27].
Figure 2: RanDumb projects the datapoints to a high-dimensional space to create a clearer separation between classes. Subsequently, it corrects the anisotropy across feature dimensions, scaling them to be unit variance each. This allows cosine similarity to accurately separates classes. The figure is adapted from [48].
@@
Key Properties. RanDumb needs no storage of exemplars and requires only one pass over the data in a one-sample-per-timestep fashion. Furthermore, it only requires online estimation of the sample covariance matrix and nearest class mean.
@@
Key Finding 1: Poor Representation Learning. We compare RanDumb with leading methods: VAE-GC [63] in Table 1 (left, middle) and SLCA [78] in Table 1 (right, middle). The primary distinction between them is their representation: RanDumb uses a fixed function (random/pretrained network), whereas VAE-GC and SLCA further continually trained deep networks. RanDumb consistently surpasses VAE-GC and SLCA by wide margins of 5-15%. This shows that state-of-the-art online continual learning algorithms fail to learn effective representations across standard exemplar-free continual learning benchmarks.
Finding 2: Over-Constrained Benchmarks. Given the demonstrated limitations of existing continual representation learning methods, an important question arises: Can better methods learn more effective representations? To explore this, we evaluated the performance of RanDumb against joint training, models trained without continual learning constraints, in both online and offline settings, as shown in Table 1 (left, bottom) and Table 1 (right, bottom). Our straightforward baseline, RanDumb, bridges 70-90% of the performance gap relative to the respective joint classifiers in both scenarios. This significant recovery of performance by such a simple method suggests that if our goal is to advance the study of representation learning, current benchmarks may be overly restrictive and not conducive to truly effective representation learning.
We highlight that the goal in our work is not to introduce a state-of-the-art continual learning method, but challenge prevailing assumptions and open a discussion on the efficacy of representation learning in continual learning algorithms, especially in online and low-exemplar scenarios.
@@ 2 RanDumb: Mechanism & Intuitions
RanDumb has two main elements: random projection and the dumb learner. We illustrate the mechanism of RanDumb using three toy examples in Figure 2. To classify a test sample $\bf x_{\rm test}$, we start with a simple classifier, the nearest class mean (NCM). It predicts the class among $C$ classes by highest value of the similarity function $f$ among class means $\mathbf{\mu}_{i}$:
$y_{\textrm{pred}}=\operatorname*{arg\,max}_{{i}\in\{1,\ldots,|C|\}}f({\bf x_{\textrm{test}}},\mathbf{\mu}_{i}),$ (1)
$\text{where}\quad f({\bf x_{\textrm{test}}},\mathbf{\mu}_{i}):={\bf x_{\textrm{test}}}^{\top}{\mathbf{\mu}_{i}}$ (2)
and $\mathbf{\mu}_{i}$ are the class-means in the pixel space: $\mathbf{\mu}_{i}=\frac{1}{|C_{i}|}\sum_{\mathbf{x}\in C_{i}}\mathbf{x}$. RanDumb adds two additional components to this classifier: 1) Kernelization and 2) Decorrelation.
Kernelization: Classes are typically not linearly separable in the pixel space, unlike in the feature space of deep models. Hence, we apply the kernel trick to embed the pixels in a better representation space, computing all distances between the data and class-means in this embedding space. This phenomena is illustrated on three toy examples to build intuitions in Figure 1 (right, Embed). We use an RBF-Kernel, which for two points $\mathbf{x}$ and $\mathbf{y}$ is defined as: $K_{\text{RBF}}(\mathbf{x},\mathbf{y})=\exp\left(-\gamma\|\mathbf{x}-\mathbf{y}\|^{2}\right)$ where $\gamma$ is a scaling parameter. However, calculating the RBF kernel is not possible due to the online continual learning constraints preventing computation of pairwise-distance between all points. Hence, we use a data-independent approximation, random Fourier projection $\phi(\mathbf{x})$, as given in [52]:
@@
We keep these $\omega$ bases fixed throughout online learning. Thus, our modified similarity function from Equation 1 is:
@@
$\bar{\mathbf{\mu}}_{i}=\frac{1}{|C_{i}|}\sum_{\mathbf{x}\in C_{i}}\phi(\mathbf{x})$
Decorrelation: Projected raw pixels have feature dimensions with different variances (anisotropic). Hence, instead of naively computing ${\bf\phi(x_{\rm test})}^{\top}{\bar{\mathbf{\mu}}_{i}}$, we further decorrelate the feature dimensions using a Mahalonobis distance with the shrinked covariance matrix $\mathbf{S}$ using OAS shrinkage [15], inverse obtained by least squares minimization ($\mathbf{S}+\lambda\mathbf{I}$). We illustrate this phenomena as well on three toy examples in Figure 1 (right, Decorrelate) to build intuitions. Our similarity function finally is:
@@
Online Computation. Our random projection is fixed before seeing any data. During continual learning, we only perform online update on the running class mean and empirical covariance matrix [^1] Online update for the inverse of the covariance matrix is possible using the Sherman–Morrison formula..
3 Experiments
We compare RanDumb with algorithms across online continual learning benchmarks with an emphasis on exemplar-free and low-exemplar storage regime.
Table 2: Overview of Benchmarks
@@ 3 Experiments
Benchmarks. The benchmarks which we used in our experiments are summarized in Table 2. We aim for a comprehensive coverage and show results on four standard online continual learning benchmarks (A, B, D, E) which reflect the latest trends (’22-’24) across exemplar-free, contrastive-training [^2] Benchmark B is split into two sections: (B1) methods that do not rely on contrastive learning and heavy augmentation, and (B2) approaches that incorporate contrastive learning and extra augmentations., meta-continual learning, and network-expansion based approaches respectively. We also evaluate on a rehearsal-free offline continual learning benchmark C. These benchmarks are ordered by increasingly relaxed constraints, moving further away from the training scenario of RanDumb. Benchmark A closely matches RanDumb with one class per timestep and no stored exemplars. Benchmark B, D, E progressively relax the constraints on exemplars and classes per timestep. Benchmark C and E remove the online constraint by allowing unrestricted training and sample access within a task without exemplar-storage of past tasks.
We further test on exemplar-free scenarios in offline continual learning using Benchmark F [67] with the challenging one-class per task constraint borrowed from [75], i.e. testing over longer timespans. This benchmark allows using pretrained models along with unrestricted training time and access to all class samples at each timestep. However, RanDumb is restricted to learning from a single pass seeing only one sample at a time. RanDumb only learns a linear classifier over a given pretrained model in Benchmark F.
We use LAMDA-PILOT [61] codebase for all methods, except RanPAC and SLDA for which use their codebases. We use the original hyperparameters. We only change initial classes to 2 and number of classes per task to 1 and test using both ImageNet21K & ImageNet1K ViT-B/16 models.
Table 3: Benchmark A (Ref: Table 1 from PEC [75]). We compare RanDumb in a 1-class per task setting referred as ‘Dataset (num_tasks/1)’. We observe that RanDumb outperforms all approaches across all datasets by 2-6% margins, with an exception of latest work PEC [75] on CIFAR10.
Method Memory MNIST CIFAR-10 CIFAR-100 miniImageNet
(10/1) (10/1) (100/1) (100/1)
@@
Rehearsal Based Methods ER [13] 500 84.4$\pm$ 0.3 40.6$\pm$ 1.1 12.5$\pm$ 0.3 5.7$\pm$ 0.2
A-GEM [12] 500 59.8$\pm$ 0.8 10.2$\pm$ 0.1 1.0$\pm$ 0.0 1.1$\pm$ 0.1
iCaRL [54] 500 83.1$\pm$ 0.3 37.8$\pm$ 0.4 5.7$\pm$ 0.1 7.5$\pm$ 0.1
BiC [70] 500 86.0$\pm$ 0.4 35.9$\pm$ 0.4 6.4$\pm$ 0.3 1.5$\pm$ 0.1
ER-ACE [10] 500 87.8$\pm$0.2 39.9$\pm$0.5 8.2$\pm$0.2 5.7$\pm$0.2
DER [9] 500 91.7$\pm$ 0.1 40.0$\pm$ 1.5 1.0$\pm$ 0.1 1.0$\pm$ 0.0
DER++ [9] 500 91.9$\pm$ 0.2 35.6$\pm$ 2.4 6.2$\pm$ 0.4 1.4$\pm$ 0.1
X-DER [8] 500 83.0$\pm$ 0.1 43.2$\pm$ 0.5 15.6$\pm$ 0.1 8.2$\pm$ 0.4
GDumb [49] 500 91.0$\pm$0.2 50.7$\pm$0.7 8.2$\pm$0.2 -
Rehearsal Free Methods EWC [33] 0 10.1$\pm$ 0.0 10.6$\pm$ 0.4 1.0$\pm$ 0.0 1.0$\pm$ 0.0
SI [76]) 0 12.7$\pm$ 1.0 10.1$\pm$ 0.1 1.1$\pm$0.0 1.0$\pm$0.1
LwF [37] 0 11.8 $\pm$ 0.6 10.1$\pm$ 0.1 0.9$\pm$0.0 1.0$\pm$ 0.0
LT [77] 0 10.9$\pm$ 0.9 10.0$\pm$ 0.2 1.1$\pm$ 0.1 1.0$\pm$ 0.0
Gen-NCM [31] 0 82.0$\pm$ 0.0 27.7$\pm$ 0.0 10.0$\pm$ 0.0 7.5$\pm$ 0.0
Gen-SLDA [27] 0 88.0$\pm$ 0.0 41.4$\pm$ 0.0 18.8$\pm$ 0.0 12.9$\pm$ 0.0
VAE-GC [63] 0 84.0$\pm$ 0.5 42.7$\pm$ 1.3 19.7$\pm$ 0.1 12.1$\pm$ 0.1
PEC [75] 0 92.3$\pm$ 0.1 58.9$\pm$ 0.1 26.5$\pm$ 0.1 14.9$\pm$ 0.1
RanDumb (Ours) 0 98.3 (+5.9) 55.6 (-3.3) 28.6 (+2.1) 17.7 (+2.8)
Table 4: Benchmark B.1 (Ref: Table adopted from OnPro [68], OCM[25]) We compare RanDumb in many-classes per task setting referred as ‘Dataset (num_tasks/num_classes_per_task)’. We categorize memory buffer sizes with ‘M’. RanDumb outperforms the competing approaches without heavy-augmentations by 3-20% margins despite being exemplar free. Only in one case, it is second best.
Method MNIST CIFAR10 CIFAR100 CIFAR100 TinyImageNet
(5/2) (5/2) (10/10) (50/2) (100/2)
$M=0.1k$ $M=0.1k$ $M=0.2k$ $M=0.5k$ $M=1k$ $M=1k$ $M=1k$ $M=2k$
AGEM [12] 56.9±5.2 17.7±0.3 22.7±1.8 5.8±0.2 5.9±0.1 1.8±0.2 0.8±0.1 0.9±0.1
GSS [4] 70.4±1.5 18.4±0.2 26.9±1.2 8.1±0.2 11.1±0.2 4.3±0.2 1.1±0.1 3.3±0.5
ER [13] 78.7±0.4 19.4±0.6 29.7±1.0 8.7±0.3 15.7±0.3 8.3±0.3 1.2±0.1 5.6±0.5
ASER [58] 61.6±2.1 20.0±1.0 27.8±1.0 11.0±0.3 16.4±0.3 9.6±1.3 2.2±0.1 5.3±0.3
MIR [3] 79.0±0.5 20.7±0.7 37.3±0.3 9.7±0.3 15.7±0.2 12.7±0.3 1.4±0.1 6.1±0.5
ER-AML [10] 76.5±0.1 - 40.5±0.7 - 16.1±0.4 - - 5.4±0.2
iCaRL [54] - 31.0±1.2 33.9±0.9 12.8±0.4 16.5±0.4 - 5.0±0.3 6.6±0.4
DER++ [9] 74.4±1.1 31.5±2.9 44.2±1.1 16.0±0.6 21.4±0.9 9.3±0.3 3.7±0.4 5.1±0.8
GDumb [49] 81.2±0.5 23.3±1.3 35.9±1.1 8.2±0.2 18.1±0.3 18.1±0.3 4.6±0.3 12.6±0.1
CoPE [19] - 33.5±3.2 37.3±2.2 11.6±0.4 14.6±1.3 - 2.1±0.3 2.3±0.4
DVC [25] - 35.2±1.7 41.6±2.7 15.4±0.3 20.3±1.0 - 4.9±0.6 7.5±0.5
Co²L [11] 83.1±0.1 - 42.1±1.2 - 17.1±0.4 - - 10.1±0.2
R-RT [6] 89.1±0.3 - 45.2±0.4 - 15.4±0.3 - - 6.6±0.3
CCIL [46] 86.4±0.1 - 50.5±0.2 - 18.5±0.3 - - 5.6±0.9
IL2A [81] 90.2±0.1 - 54.7±0.5 - 18.2±1.2 - - 5.5±0.7
BiC [70] 90.4±0.1 - 48.2±0.7 - 21.2±0.3 - - 10.2±0.9
SSIL [1] 88.2±0.1 - 49.5±0.2 - 26.0±0.1 - - 9.6±0.7